During a race on an oval track, a car travels at an average speed of 200 km/hr. (a) how far did it travel in 45.0 min? (b) determine its average velocity at the end of its third lap


1
Expert's answer
2021-03-29T08:58:53-0400

(a)

"s=vt=200\ \dfrac{km}{h}\cdot45\ min\cdot\dfrac{1\ h}{60\ min}=150\ km."

(b) The average velocity can be found as follows:


"v_{avg}=\dfrac{Displacement}{Time\ taken}."

Since the displacement of the car equals zero (the car returns to the point from which it was start), the average velocity equals zero.